# A Stitch in Time: Learning Procedural Workflow via Self-Supervised Plackett–Luce Ranking

### v1

@@ Abstract
Procedural activities, ranging from routine cooking to complex surgical operations, are highly structured as a set of actions conducted in a specific temporal order. Despite their success on static images and short clips, current self-supervised learning methods often overlook the procedural nature that underpins such activities. We expose the lack of procedural awareness in current SSL methods with a motivating experiment: models pretrained on forward and time-reversed sequences produce highly similar features, confirming that their representations are blind to the underlying procedural order. To address this shortcoming, we propose PL-Stitch, a self-supervised framework that harnesses the inherent temporal order of video frames as a powerful supervisory signal. Our approach integrates two novel probabilistic objectives based on the Plackett-Luce (PL) model. The primary PL objective trains the model to sort sampled frames chronologically, compelling it to learn the global workflow progression. The secondary objective, a spatio-temporal jigsaw loss, complements the learning by capturing fine-grained, cross-frame object correlations. Our approach consistently achieves superior performance across five surgical and cooking benchmarks. Specifically, PL-Stitch yields significant gains in surgical phase recognition (e.g., +11.4 pp k-NN accuracy on Cholec80) and cooking action segmentation (e.g., +5.7 pp linear probing accuracy on Breakfast), demonstrating its effectiveness for procedural video representation learning.
@@
Figure 1: Procedural awareness in self-supervised learning (SSL). (a) Prior SSL methods are procedural-agnostic, learning features that fail to capture the procedural structure. (b) We pretrain PL-Stitch and prior SSL methods on the Breakfast dataset [Kuehne2014TheActivities] using forward and backward sequences. The violin plot shows the cosine distance between the feature vectors of the same video frames. Baselines exhibit procedural agnosticism (low distance), whereas PL-Stitch’s high distance validates its procedural awareness.
@@ 1 Introduction
To provide insights on this question, we conducted an experiment that motivates our work. We pretrained leading SSL models on both forward (chronological) and backward (time-reversed) video sequences of the Breakfast dataset [Kuehne2014TheActivities]. At inference, we feed the same video frames to both the forward and backward trained encoders and compute the cosine distance between their resulting features. As shown in Fig. 1, these models produce nearly identical features regardless of temporal direction, confirming their representations are blind to the underlying procedural order. This demonstrates that while they can recognize an action (e.g., grinding beans), they fail to capture its essential temporal context (knowing it must come before brewing).
@@
This critical failure stems from the inherent design of prevalent SSL objectives, which are centered on local tasks like instance discrimination [Caron2021EmergingTransformers, Chen2020ARepresentations, Chen2020ImprovedLearning] or masked reconstruction [He2022MaskedLearners, HangboBaoandLiDongandSonghaoPiaoandFuruWei2022BEIT:TRANSFORMERS, Zhou2022IBOT:Tokenizer, Wang2023VideoMAEMasking]. Consequently, despite achieving strong performance in static image analysis [Caron2021EmergingTransformers, Chen2020ImprovedLearning, Chen2020ARepresentations, He2022MaskedLearners, Zhou2022IBOT:Tokenizer] and short atomic clip modeling [Feichtenhofer2021ALearning, Wang2022BEVT:Transformers, ZhanTongandYibingSongandJueWangandLiminWang2022VideoMAE:Pre-Training, Wang2023VideoMAEMasking], current SSL approaches remain fundamentally procedural-agnostic (Fig. 1(a)). They learn robust features for what is in a frame, but not when that frame occurs in the sequence.
@@
The video branch serves as our main temporal ranking objective. It learns global workflow progression by tasking the model to predict the correct chronological order of sampled frames. Importantly, we formulate this as a listwise ranking problem using the probabilistic Plackett-Luce (PL) model [Plackett1975ThePermutations, Luce1977TheYears]. Our approach has key advantages over traditional temporal ordering tasks: 1) Its listwise formulation is more efficient and globally consistent than suboptimal, pairwise comparisons [Hu2021ContrastLearning, Lee2017UnsupervisedSequences, Sermanet2017Time-ContrastiveObservation], as it optimizes the ordering of a sequence of $k$ elements in a single step, providing a global signal rather than relying on $\mathcal{O}(k^{2})$ fragmented local comparisons. 2) Its probabilistic, ranking-based nature is a more robust fit for modeling relative order than classification tasks [Misra2016ShuffleVerification, Wei2018LearningTime, Xu2019Self-SupervisedPrediction, Fernando2017Self-SupervisedNetworks]. Permutation classification tasks mistreat this relative problem by using absolute labels, penalizing near-correct orderings (minor sorting mistakes) as fully wrong. In contrast, our PL-based ranking model computes a probability distribution over permutations, allowing penalties to scale with error severity.
The complementary image branch learns fine-grained, local features by jointly optimizing two objectives: 1) our novel spatio-temporal jigsaw objective, which uses adjacent past and future frames as temporal context to learn object correspondence, and 2) a masked image modeling (MIM) loss [Zhou2022IBOT:Tokenizer] for robust semantic representations.
@@ 2 Related Work
Self-supervised Visual Representation Learning. A widely adopted approach in self-supervised visual representation learning is contrastive learning, which seeks to learn an embedding space that groups similar images together. Pioneer methods like MoCo [He2020MomentumLearning, Chen2020ImprovedLearning, Chen2021AnTransformers] and SimCLR [Chen2020ARepresentations] learn this space by pulling augmented positive views of an image closer while pushing negative views apart. To prevent representational collapse without negative pairs, self-distillation methods like DINO [Caron2021EmergingTransformers] employ a student-teacher framework to match augmented views. Furthermore, these contrastive and distillation approaches are also adapted to the video domain by leveraging temporal dynamics [Feichtenhofer2021ALearning, Recasens2021BroadenLearning, Qian2021SpatiotemporalLearning, Sun2023MaskedLearning, Qing2023Self-SupervisedConsistency, Wang2021EnhancingMotion]. More recently, Masked Image Modeling (MIM) has become prominent, tasking a model with reconstructing masked patches at the pixel level (MAE [He2022MaskedLearners]) or latent-space level (BEiT [HangboBaoandLiDongandSonghaoPiaoandFuruWei2022BEIT:TRANSFORMERS], iBOT [Zhou2022IBOT:Tokenizer]). Building on the iBOT objective, DINOv2 [Oquab2023DINOv2:Supervision] and DINOv3 [Simeoni2025DINOv3] later scaled this approach with massive data to boost performance. Masked Video Modeling (MVM) [Wang2023VideoMAEMasking, ZhanTongandYibingSongandJueWangandLiminWang2022VideoMAE:Pre-Training, Pei2024VideoMAC:ConvNets, Wu2023DropMAE:Tasks, Feichtenhofer2022MaskedLearners, Wang2022BEVT:Transformers, Huang2023MGMAE:Autoencoding] extends this reconstruction paradigm to video. Methods like VideoMAE [ZhanTongandYibingSongandJueWangandLiminWang2022VideoMAE:Pre-Training] learn representations by reconstructing randomly masked spatio-temporal tubes. However, a key limitation is its symmetric treatment of space and time [Feichtenhofer2022MaskedLearners, Huang2023MGMAE:Autoencoding], which ignores the inherent causal progression of the temporal dimension and makes such methods suboptimal for learning the dynamics of procedural events.
@@
Temporal Correspondence Learning. Understanding temporal correspondence between frames is crucial for modeling procedural activities. Early supervised methods using optical flow and motion estimation were effective but required costly pixel-level annotations [Teed2020RAFT:Flow, Sun2018PWC-Net:Volume, Xu2017AccurateProcessing, Ilg2017FlowNetNetworks]. To overcome this issue, self-supervised approaches learned spatio-temporal correspondence by tracking objects or patches across frames [Wang2019LearningTime, Zhang2023BoostingLearning, Jabri2020Space-timeWalk, Xu2021RethinkingPerspective, Li2019Joint-taskCorrespondence]. More recently, temporal predictive reconstruction has emerged, where SiamMAE [Gupta2023SiameseAutoencoders] predicts a future frame from the past and T-CoRe [Liu2025WhenLearning] using both past and future frames to reconstruct a central frame. While these methods excel at learning fine-grained correspondence, they lack a mechanism to capture video’s long-term procedural structure. Training models to order frame sequences has also been proposed [Xu2019Self-SupervisedPrediction, Wei2018LearningTime, Lee2017UnsupervisedSequences, Fernando2017Self-SupervisedNetworks, Misra2016ShuffleVerification, Hu2021ContrastLearning, Sermanet2017Time-ContrastiveObservation]. However, this approach is limited by suboptimal objectives, such as pairwise comparisons that provide a local signal [Hu2021ContrastLearning, Lee2017UnsupervisedSequences, Sermanet2017Time-ContrastiveObservation], or permutation classification tasks that miscast relative ordering as absolute classification [Xu2019Self-SupervisedPrediction, Wei2018LearningTime, Misra2016ShuffleVerification]. Our work, PL-Stitch, addresses this gap by reformulating temporal understanding as a more robust, probabilistic listwise ranking problem, which we argue is a more natural and direct way to model sequences.
@@
Procedural Activity Understanding. Modeling procedural activities involves recognizing both the individual action steps and their temporal sequence. Many recent works rely on densely annotated videos to localize and predict the order of steps [Chang2020ProcedureVideos, Kuehne2014TheActivities, Tang2019COIN:Analysis, Zhukov2019Cross-TaskVideos, Damen2021TheBaselines]. To reduce the reliance on expensive labels, subsequent research has explored weakly supervised approaches. These methods typically operate on the premise that language provides semantic context for understanding complex procedures, leveraging transcripts [Bojanowski2015Weakly-SupervisedText, Zhao2022PSupervision, Zhukov2019Cross-TaskVideos], video–narration alignment [Miech2020End-to-EndVideos, Elhamifar2020Self-supervisedVideos, Zhong2023LearningNarrations, Yuan2024Procedure-awareAugmentation], text-derived procedural graphs [Zhou2023Procedure-AwareUnderstanding], or text-guided step discovery [Shah2023STEPs:Videos]. In contrast, we show that PL-Stitch can learn robust procedural representations solely from the video signal in a fully self-supervised regime, effectively eliminating the need for auxiliary modalities.
@@ 3.1 Problem Formulation
Endo-FM [Wang2023FoundationPre-train] S Private data ViT-B/16 $51.5$ $43.1$ $46.8$ $62.7$ $53.9$ $55.9$ $53.8$ $49.8$ $44.8$
EndoViT [Batic2024EndoViT:Images] Merged public data ViT-B/16 $45.4$ $38.7$ $40.2$ $46.9$ $39.0$ $45.8$ $38.5$ $37.2$ $34.6$
LemonFM [Che2025LEMON:Settings] LEMON ConvNeXt-B $74.7$ $64.5$ $73.1$ $73.9$ $65.8$ $69.5$ $68.4$ $62.4$ $65.8$
MAE [He2022MaskedLearners] G LEMON ViT-B/16 $35.5$ $32.0$ $35.9$ $54.9$ $43.4$ $47.2$ $40.7$ $38.1$ $36.4$
VideoMAEv2 [Wang2023VideoMAEMasking] LEMON ViT-B/16 $49.8$ $42.4$ $46.8$ $55.8$ $48.5$ $50.0$ $50.7$ $41.3$ $45.8$
DINO [Caron2021EmergingTransformers] LEMON ViT-B/16 $74.9$ $65.0$ $72.8$ $72.2$ $67.1$ $69.4$ $67.6$ $62.8$ $65.2$
iBOT [Zhou2022IBOT:Tokenizer] LEMON ViT-B/16 $76.3$ $65.1$ $75.3$ $74.6$ $67.6$ $70.3$ $71.0$ $64.5$ $68.0$
T-CoRe [Liu2025WhenLearning] LEMON ViT-B/16 $67.9$ $58.0$ $61.8$ $65.1$ $59.4$ $62.3$ $61.8$ $57.1$ $60.2$
@@ 3.2 Overview
Our goal is to learn a self-supervised representation capable of capturing both procedural temporal correspondence and fine-grained spatio-temporal relationships. To that end, we propose PL-Stitch, a multi-task framework that jointly trains a shared backbone encoder, $f_{\theta}$, using two complementary branches as illustrated in Fig. 3. The Video branch (Sec. 3.4) is designed to learn temporal ordering across variable time scales by training the encoder to predict the correct chronological sequence of a sampled clip ($\mathcal{L}_{\text{vid}}$). In parallel, the Image branch (Sec. 3.5) learns robust local features through two tasks: our novel spatio-temporal jigsaw ($\mathcal{L}_{\text{jigsaw}}$), which learns fine-grained correspondence using adjacent frames as context and a masked image modeling objective ($\mathcal{L}_{\text{MIM}}$) based on iBOT [Zhou2022IBOT:Tokenizer]. Importantly, both the temporal ordering and the spatio-temporal jigsaw tasks are framed as listwise ranking problems, optimized with a unified Plackett-Luce objective (Sec. 3.3). The entire training procedure is summarized in Algorithm 1.
@@ 3.3 Plackett-Luce Ranking Formulation
Both our temporal and spatial learning tasks are framed as a listwise ranking problem. We address this using the Plackett-Luce (PL) distribution [Plackett1975ThePermutations, Luce1977TheYears], a probabilistic framework for defining a distribution over all $K!$ possible permutations of $K$ items. The PL model is derived from Luce’s Choice Axiom [Luce1977TheYears] (independence from irrelevant alternatives), which states that the odds of choosing item $i$ over item $j$ depend only on those two items. This assumption aligns well with relative time ordering, where the preference between two frames is invariant to the rest of the sequence. Given a set of $K$ items, the PL model is parameterized by a vector of real-valued scores $s\in\mathbb{R}^{K}$. The probability of observing a specific permutation $r$ of the set $\{1,\dots,K\}$ is given by the likelihood:
@@ 3.4 Video Branch: Listwise Temporal Ranking
Table 2: Comparison with state-of-the-art self-supervised methods on cooking datasets. We use the egocentric cooking dataset GTEA [Fathi2011LearningActivities] and the third-person cooking dataset Breakfast [Kuehne2014TheActivities]. All methods are evaluated on a ViT-B/16 backbone using the linear probing and $k$-NN protocols.
@@
VideoMAEv2 [Wang2023VideoMAEMasking] $30.2$ $22.8$ 23.523.517.817.810.010.0 $26.3$
DINO [Caron2021EmergingTransformers] $52.2$ $53.9$ 55.355.346.746.728.328.3 $60.0$
iBOT [Zhou2022IBOT:Tokenizer] $51.3$ $57.2$ 59.359.349.049.027.427.4 $59.2$
T-CoRe [Liu2025WhenLearning] $47.2$ $50.0$ 49.349.342.842.820.920.9 $42.7$
@@
iBOT [Zhou2022IBOT:Tokenizer] $15.7$ $12.0$ 7.27.24.84.82.62.6 $7.5$
T-CoRe [Liu2025WhenLearning] $12.0$ $12.5$ 6.16.14.04.01.81.8 $6.1$

### v2

@@ Abstract
Procedural activities, ranging from routine cooking to complex surgical operations, are highly structured sequences of actions performed in a specific temporal order. Despite the success of current self-supervised learning (SSL) methods on static images and short clips, these models often overlook the underlying sequential structure of such activities. We expose this lack of procedural awareness with a motivating experiment: models pretrained on forward and time-reversed sequences produce highly similar features, confirming that their representations are blind to the underlying procedural order. To address this shortcoming, we propose PL-Stitch, a self-supervised framework that harnesses the inherent temporal order of video frames as a powerful supervisory signal. Our approach integrates two novel probabilistic objectives based on the Plackett-Luce (PL) model. The primary PL objective trains the model to sort sampled frames chronologically, compelling it to learn the global workflow progression. The secondary objective, a spatio-temporal jigsaw loss, complements the learning by capturing fine-grained, cross-frame object correspondences. Our approach consistently achieves superior performance across five surgical and cooking benchmarks. Specifically, PL-Stitch yields significant gains in surgical phase recognition (e.g., +11.4 pp in k-NN accuracy on Cholec80) and cooking action segmentation (e.g., +5.7 pp in linear probing accuracy on Breakfast), demonstrating its effectiveness for procedural video representation learning. Code and models are available at https://github.com/visurg-ai/PL-Stitch.
@@
Figure 1: Procedural awareness in self-supervised learning (SSL). (a) Prior SSL methods are procedural-agnostic, learning features that fail to capture the procedural structure. (b) We pretrain PL-Stitch and prior SSL methods on the Breakfast dataset [24] using forward and backward sequences. The violin plot shows the cosine distance between the feature vectors of the same video frames. Baselines exhibit procedural agnosticism (low distance), whereas PL-Stitch’s high distance validates its procedural awareness.
@@ 1 Introduction
To provide insights on this question, we conducted an experiment that motivates our work. We pretrained leading SSL models on both forward (chronological) and backward (time-reversed) video sequences of the Breakfast dataset [24]. At inference, we feed the same video frames to both the forward and backward trained encoders and compute the cosine distance between their resulting features. As shown in Fig. LABEL:fig:figure1, these models produce nearly identical features regardless of temporal direction, confirming their representations are blind to the underlying procedural order. This demonstrates that while they can recognize an action (e.g., grinding beans), they fail to capture its essential temporal context (knowing it must come before brewing).
@@
This critical failure stems from the inherent design of prevalent SSL objectives, which are centered on local tasks like instance discrimination [3, 7, 8] or masked reconstruction [18, 17, 64, 50]. Consequently, despite achieving strong performance in static image analysis [3, 8, 7, 18, 64] and short atomic clip modeling [14, 51, 61, 50], current SSL approaches remain fundamentally procedural-agnostic (Fig. LABEL:fig:figure1a). They learn robust features for what is in a frame, but not when that frame occurs in the sequence.
@@
The video branch serves as our main temporal ranking objective. It learns global workflow progression by tasking the model to predict the correct chronological order of sampled frames. Importantly, we formulate this as a listwise ranking problem using the probabilistic Plackett-Luce (PL) model [35, 30]. Our approach has key advantages over traditional temporal ordering tasks: 1) Its listwise formulation is more efficient and globally consistent than suboptimal, pairwise comparisons [20, 25, 39], as it optimizes the ordering of a sequence of $k$ elements in a single step, providing a global signal rather than relying on $\mathcal{O}(k^{2})$ fragmented local comparisons. 2) Its probabilistic, ranking-based nature is a more robust fit for modeling relative order than classification tasks [31, 55, 58, 15]. Permutation classification tasks mistreat this relative problem by using absolute labels, penalizing near-correct orderings (minor sorting mistakes) as fully wrong. In contrast, our PL-based ranking model computes a probability distribution over permutations, allowing penalties to scale with error severity.
The complementary image branch learns fine-grained, local features by jointly optimizing two objectives: 1) our novel spatio-temporal jigsaw objective, which uses adjacent past and future frames as temporal context to learn object correspondence, and 2) a masked image modeling (MIM) loss [64] for robust semantic representations.
@@ 2 Related Work
Self-supervised Visual Representation Learning. A widely adopted approach in self-supervised visual representation learning is contrastive learning, which seeks to learn an embedding space that groups similar images together. Pioneer methods like MoCo [19, 8, 9] and SimCLR [7] learn this space by pulling augmented positive views of an image closer while pushing negative views apart. To prevent representational collapse without negative pairs, self-distillation methods like DINO [3] employ a student-teacher framework to match augmented views. Furthermore, these contrastive and distillation approaches are also adapted to the video domain by leveraging temporal dynamics [14, 38, 36, 43, 37, 49]. More recently, Masked Image Modeling (MIM) has become prominent, tasking a model with reconstructing masked patches at the pixel level (MAE [18]) or latent-space level (BEiT [17], iBOT [64]). Building on the iBOT objective, DINOv2 [32] and DINOv3 [40] later scaled this approach with massive data to boost performance. Masked reconstruction paradigms have also been successfully extended from images to higher-dimensional modalities [50, 61, 34, 44, 56, 13, 51, 21], particularly videos. Methods like VideoMAE [61] learn representations by reconstructing randomly masked spatio-temporal tubes from video sequences. However, a key limitation is its symmetric treatment of space and time [13, 21], which ignores the inherent causal progression of the temporal dimension and makes such methods suboptimal for learning the dynamics of procedural events.
@@
Temporal Correspondence Learning. Understanding temporal correspondence between frames is crucial for modeling procedural activities. Early supervised methods using optical flow and motion estimation were effective but required costly pixel-level annotations [46, 42, 59, 22]. To overcome this issue, self-supervised approaches learned spatio-temporal correspondence by tracking objects or patches across frames [52, 62, 23, 60, 27]. More recently, temporal predictive reconstruction has emerged, where SiamMAE [16] predicts a future frame from the past and T-CoRe [29] using both past and future frames to reconstruct a central frame. While these methods excel at learning fine-grained correspondence, they lack a mechanism to capture video’s long-term procedural structure. Training models to order frame sequences has also been proposed [58, 55, 25, 15, 31, 20, 39]. However, this approach is limited by suboptimal objectives, such as pairwise comparisons that provide a local signal [20, 25, 39], or permutation classification tasks that miscast relative ordering as absolute classification [58, 55, 31]. Our work, PL-Stitch, addresses this gap by reformulating temporal understanding as a more robust, probabilistic listwise ranking problem, which we argue is a more natural and direct way to model sequences.
@@
Procedural Activity Understanding. Modeling procedural activities, which involves both recognizing action steps and their temporal sequence, is a long-standing challenge in video understanding. Most recent works consider training on densely-annotated videos to localize and predict the order of steps and the logical dependencies [57, 5, 24, 45, 65, 10]. To reduce this reliance on expensive labels, other methods have explored weakly-supervised approaches, such as learning from only a transcript of action steps without timestamps [2, 63, 65]. In contrast, our work operates in a fully self-supervised regime, learning a representation from the video signal alone that can be leveraged for procedural video understanding.
@@ 3.1 Problem Formulation
Endo-FM [53] S Private data ViT-B/16 $51.5$ $43.1$ $46.8$ $62.7$ $53.9$ $55.9$ $53.8$ $49.8$ $44.8$
EndoViT [1] Merged public data ViT-B/16 $45.4$ $38.7$ $40.2$ $46.9$ $39.0$ $45.8$ $38.5$ $37.2$ $34.6$
LemonFM [6] LEMON ConvNeXt-B $74.7$ $64.5$ $73.1$ $73.9$ $65.8$ $69.5$ $68.4$ $62.4$ $65.8$
MAE [18] G LEMON ViT-B/16 $35.5$ $32.0$ $35.9$ $54.9$ $43.4$ $47.2$ $40.7$ $38.1$ $36.4$
VideoMAEv2 [50] LEMON ViT-B/16 $49.8$ $42.4$ $46.8$ $55.8$ $48.5$ $50.0$ $50.7$ $41.3$ $45.8$
DINO [3] LEMON ViT-B/16 $74.9$ $65.0$ $72.8$ $72.2$ $67.1$ $69.4$ $67.6$ $62.8$ $65.2$
iBOT [64] LEMON ViT-B/16 $76.3$ $65.1$ $75.3$ $74.6$ $67.6$ $70.3$ $71.0$ $64.5$ $68.0$
T-CoRe [29] LEMON ViT-B/16 $67.9$ $58.0$ $61.8$ $65.1$ $59.4$ $62.3$ $61.8$ $57.1$ $60.2$
@@ 3.2 Overview
Our goal is to learn a self-supervised representation capable of capturing both procedural temporal correspondence and fine-grained spatio-temporal relationships. To that end, we propose PL-Stitch, a multi-task framework that jointly trains a shared backbone encoder, $f_{\theta}$, using two complementary branches as illustrated in Fig. 3. The Video branch (Sec. 3.4) is designed to learn temporal ordering across variable time scales by training the encoder to predict the correct chronological sequence of a sampled clip ($\mathcal{L}_{\text{vid}}$). In parallel, the Image branch (Sec. 3.5) learns robust local features through two tasks: our novel spatio-temporal jigsaw ($\mathcal{L}_{\text{jigsaw}}$), which learns fine-grained correspondence using adjacent frames as context and a masked image modeling objective ($\mathcal{L}_{\text{MIM}}$) based on iBOT [64]. Importantly, both the temporal ordering and the spatio-temporal jigsaw tasks are framed as listwise ranking problems, optimized with a unified Plackett-Luce objective (Sec. 3.3). The entire training procedure is summarized in Algorithm 1.
@@ 3.3 Plackett-Luce Ranking Formulation
Both our temporal and spatial learning tasks are framed as a listwise ranking problem. We address this using the Plackett-Luce (PL) distribution [35, 30], a probabilistic framework for defining a distribution over all $K!$ possible permutations of $K$ items. The PL model is derived from Luce’s Choice Axiom [30] (independence from irrelevant alternatives), which states that the odds of choosing item $i$ over item $j$ depend only on those two items. This assumption aligns well with relative time ordering, where the preference between two frames is invariant to the rest of the sequence. Given a set of $K$ items, the PL model is parameterized by a vector of real-valued scores $s\in\mathbb{R}^{K}$. The probability of observing a specific permutation $r$ of the set $\{1,\dots,K\}$ is given by the likelihood:
@@ 3.4 Video Branch: Listwise Temporal Ranking
Table 2: Comparison with state-of-the-art self-supervised methods on cooking datasets. We use the egocentric cooking dataset GTEA [12] and the third-person cooking dataset Breakfast [24]. All methods are evaluated on a ViT-B/16 backbone using the linear probing and $k$-NN protocols.
@@
VideoMAEv2 [50] $30.2$ $22.8$ 23.523.517.817.810.010.0 $26.3$
DINO [3] $52.2$ $53.9$ 55.355.346.746.728.328.3 $60.0$
iBOT [64] $51.3$ $57.2$ 59.359.349.049.027.427.4 $59.2$
T-CoRe [29] $47.2$ $50.0$ 49.349.342.842.820.920.9 $42.7$
@@
iBOT [64] $15.7$ $12.0$ 7.27.24.84.82.62.6 $7.5$
T-CoRe [29] $12.0$ $12.5$ 6.16.14.04.01.81.8 $6.1$